OpenStudy (anonymous):

y''' - y' = 25cos(2x) @TuringTest so what about that we have more than a λ. how can i use these

OpenStudy (anonymous):

i'm not entirely sure what you're asking in this case lambda usually comes with eigenvalues and functions

OpenStudy (turingtest):

for both the complimentary and particular you will have to throw in some extra x's I think. How many I'm not so sure.

OpenStudy (anonymous):

if you mean what if we have more than one solution for your coplementary

OpenStudy (anonymous):

in last question you say m-3 = 0

OpenStudy (anonymous):

when you particular = one of your complementary solutions , then you must multiply by x's until you get ne that is not

OpenStudy (anonymous):

in that m(m-1)(m+1)=0

OpenStudy (anonymous):

yes treat the primes as exponents so now you have \[m^3-m=0\] \[m(m^2-1)=0\] m=0 of multiplicity , m=1 k=1, m=-1 k =1

OpenStudy (anonymous):

writing this when you have m=0 it's just a constant and then adding x per multiplicity and for +1 you have two rational values your complementary will be \[y_c=c_1+c_2e^x+c_3e^{-x}\]

OpenStudy (turingtest):

so in this case nothing changes, because the guess for the particular is already linearly independent so nothing changes

OpenStudy (turingtest):

...for finding the particular I mean

OpenStudy (anonymous):

ok i get that

OpenStudy (anonymous):

wait i could not understand the relationship between k and m

OpenStudy (anonymous):

alright so now for the particular you have a solution that will take the form of 25cos(x) since when taking the derivative of cos you change from sin to cos, the possible solution is Acos(x)+Bsin(x)

OpenStudy (anonymous):

k=multiplicity m is a root

OpenStudy (turingtest):

i.e. how many times that particular root m occurs is a number k called the multiplicity of that root

OpenStudy (anonymous):

say if you had \[m^3=0\] it'd be m=0 k=3

OpenStudy (anonymous):

oh i get that too :D thanks

OpenStudy (anonymous):

so now that you know that a possible particular solution is in the form Acos(x)+Bsin(x)... just take the derivatives like turing did last time and plug them into the equation. Then solve

OpenStudy (anonymous):

@Outkast3r09 when we choose particular solution as Acos(x) + B sin (x) the equ is 0 = 25 cos x :/

OpenStudy (anonymous):

0 = 25cos (x) i am sorry

OpenStudy (anonymous):

ok i did :D

OpenStudy (anonymous):

Let me make sure hold one

OpenStudy (anonymous):

it was 25cos(2x) i calculated as 25cosx :S

OpenStudy (anonymous):

ahh yes that's my bad also your particular will be in the form Acos(2x)+Bsin(2x)

OpenStudy (anonymous):

@Outkast3r09 but in the end i found 6Asin(2x)-6Bcos(2x) = 25 cos(2x) but answer should be -5sinxcosx = -2.5sin 2x

OpenStudy (anonymous):

did i make a mistake somewhere

OpenStudy (anonymous):

\[y_p=Acos(2x)+Bsin(2x)\] \[y_p'=-2sin(2x)+2cos(2x)\] \[y_p''=-4cos(2x)+-4sin(2x)\] \[y_p'''=6sin(2x)-6cos(2x)\]

OpenStudy (anonymous):

\[6asin(2x)-6bcos(2x)-2asin(2x)+2bcos(2x)=25cos(2x)\]

OpenStudy (anonymous):

i forgot about the a's and b's but you can see how they got there correct?

OpenStudy (anonymous):

wait is not it wrong ?

OpenStudy (anonymous):

yp′′′=8sin(2x)−8cos(2x)

OpenStudy (anonymous):

ahh yes lol sorry i'm doing this all through the computer -.- \[8asin(2x)-8bcos(2x)-(-2asin(2x)+2bcos(2x)=25cos(2x)\] \[8asin(2x)-8bcos(2x)+2asin(2x)-2bcos(2x)=25cos(2x)\] \[10asin(2x)-10bcos(2x)=25cos(2x)\] -10b=25 10a=0 25/-10=b=-2.5

OpenStudy (anonymous):

10a=0 ....a=0

OpenStudy (anonymous):

so your particular is 0cos(2x)-2.5cos(2x) they just use the double angle theorem

OpenStudy (anonymous):

oh dudddeee :( i forgot the minus

OpenStudy (anonymous):

-2.5cos(2x)=-2.5(2sin(x)cos(x)=-5sin(x)cos(x)

OpenStudy (anonymous):

thank u so much :/ i am sorry i made u tired